OpenStudy (anonymous):

x^2-ax+a^2=0 find the discriminant of the equation

OpenStudy (anonymous):

\[x ^{2}-ax+a ^{2}=0 \]this is the equation correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

What does the question ask you to do?

OpenStudy (anonymous):

to find the discriminant

OpenStudy (jdoe0001):

\(\bf ax^2+bx+c=0\qquad discriminant \implies b^2-4ac\\ \quad \\ x^2-ax+a^2=0\qquad \implies \textit{what do you think?}\)

OpenStudy (jdoe0001):

notice your coefficients

OpenStudy (anonymous):

I'm following jdoe

OpenStudy (anonymous):

how do I deal with a^2

OpenStudy (jdoe0001):

well, is just a variable, thus your discriminant will be in terms of "a"

OpenStudy (anonymous):

a=1, b=a, c=a^2?

OpenStudy (jdoe0001):

\(\begin{array}{llll} &a&b&c\\ &1x^2&-ax&+a^2=0 \end{array}\qquad \bf \implies b^2-4ac\implies (-a)^2-4(1)(a^2)\)

OpenStudy (anonymous):

-3a^2

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Wow